# How does Transformer Learn Implicit Reasoning?

### v1

@@ Appendix B Dataset Details
Our data construction pipeline is adapted from the open-source code released by Wang et al. [24], with modifications to support fine-grained query-level control.
@@
First, we fix a relatively high ratio of Train-II to ID Triples (7.2:1), ensuring that compositional supervision dominates over direct fact learning. While prior work [24] has shown that a high Train-II / ID ratio is necessary for generalization on Test-II queries, we do not treat this ratio as a variable of interest. Instead, we maintain it at a sufficient level to ensure the emergence of in-distribution reasoning, and shift our focus to more challenging phenomena such as cross-distribution generalization and the underlying mechanisms that support generalizations.
@@ G.1 Training Details
Our training procedure largely follows the public implementation provided by Wang et al. [24], with a few modifications to accommodate our specific experimental setting.
@@ G.2 Scaling Analysis: Dynamics and Alignment in Larger Models
We extend our main experimental setup by training a larger model, Qwen2.5-1.5B [30], under the same base configuration. Our goal is to examine whether the developmental trajectory of multi-hop reasoning observed in smaller models persists at scale, and to further investigate the alignment between behavioral accuracy and representational metrics.
@@ Appendix H Validation of Phase III Emergence via ID/OOD Ratio Ablation
Experimental Setup. We begin with the full base configuration, which includes all atomic triples (both ID and OOD) and the complete set of Train-II queries. We fix the total number of atomic triples and vary the ID/OOD ratio while keeping all other components of training unchanged. Specifically, we test three settings: 80% ID / 20% OOD, 50% ID / 50% OOD, and 30% ID / 70% OOD. In all cases, the Train-II / ID ratio is held constant, as prior work Wang et al. [24] identifies this as a critical factor for Phase II generalization.

### v2

@@ Appendix B Dataset Details
Our data construction pipeline is adapted from the open-source code released by wang2024grokking, with modifications to support fine-grained query-level control.
@@
First, we fix a relatively high ratio of Train-II to ID Triples (7.2:1), ensuring that compositional supervision dominates over direct fact learning. While prior work [wang2024grokking] has shown that a high Train-II / ID ratio is necessary for generalization on Test-II queries, we do not treat this ratio as a variable of interest. Instead, we maintain it at a sufficient level to ensure the emergence of in-distribution reasoning, and shift our focus to more challenging phenomena such as cross-distribution generalization and the underlying mechanisms that support generalizations.
@@ G.1 Training Details
Our training procedure largely follows the public implementation provided by wang2024grokking, with a few modifications to accommodate our specific experimental setting.
@@ G.2 Scaling Analysis: Dynamics and Alignment in Larger Models
We extend our main experimental setup by training a larger model, Qwen2.5-1.5B, under the same base configuration. Our goal is to examine whether the developmental trajectory of multi-hop reasoning observed in smaller models persists at scale, and to further investigate the alignment between behavioral accuracy and representational metrics.
@@ Appendix H Validation of Phase III Emergence via ID/OOD Ratio Ablation
Experimental Setup. We begin with the full base configuration, which includes all atomic triples (both ID and OOD) and the complete set of Train-II queries. We fix the total number of atomic triples and vary the ID/OOD ratio while keeping all other components of training unchanged. Specifically, we test three settings: 80% ID / 20% OOD, 50% ID / 50% OOD, and 30% ID / 70% OOD. In all cases, the Train-II / ID ratio is held constant, as prior work wang2024grokking identifies this as a critical factor for Phase II generalization.